OpenStudy (anonymous):

solve -5/4x >with a line under it -3/8

OpenStudy (curry):

10x/3

OpenStudy (anonymous):

\[-\frac{5}{4}x\geq -\frac{3}{8}\]

OpenStudy (anonymous):

oh gimme a break!

OpenStudy (anonymous):

sleeping again... zzzzzzz

OpenStudy (saifoo.khan):

xD

OpenStudy (saifoo.khan):

NO, not this time.

OpenStudy (anonymous):

zzzzzzzzzz hic zzzzzzzzz

OpenStudy (saifoo.khan):

c'MON!

OpenStudy (saifoo.khan):

=(

OpenStudy (anonymous):

multiply both sides by \[-\frac{5}{4}\] to get x by itself and done forget to s \[\color{red}{\text{switch the inequality}}\]

OpenStudy (saifoo.khan):

final answer?

OpenStudy (anonymous):

\[x\leq \frac{3}{10}\]

OpenStudy (anonymous):

zzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzz

OpenStudy (anonymous):

i gotta watch you all night or what?

OpenStudy (saifoo.khan):

i just slept for 2 hrs! + im fasting too. lol.

OpenStudy (anonymous):

ramadan?